Prove the following theorem: Let X ~ χ, and Y ~ χ21. If X and Y are independent, then
Answer #1

For,

X sim chi^2_n

We have the Moment generating function as :

exp (ta exp(tz) z/2 1 exp(-/2)dz 2mT(n/2) 2exp((t- 22T(n/2) For the case where t<, using the change-of-variable r-(-t)z we have Mx)2(n/2) = (1-2) n/2·一一1 expdr. = (1-2t) n/2

My (t) = (1-2)-n/2

Similarly,

2

Moment generating function will be :

My (t) = (1-2t)-m/2

Now,

The moment generating function of a sum of mutually independent random variables is just the product of their moment generating functions

Moment generating function of X+Y is defined as :

M_{X+Y}(t) = M_X(t).M_Y(t)

(1-20-n/2.(1-20-rn/2 =

= (1-2t) -(n+m)/2

The above moment generating function is of Chi square distribution with degrees of freedom (n+m)

Hence, proved.